A goldsmith has two alloys of gold, the first being 70% pure and the second being 60% pure. How many ounces of the 60% pure gold must be used to make 100 ounces of an alloy which will be 66% gold?

Answer by stanbon(75887)

A goldsmith has two alloys of gold, the first being 70% pure and the second being 60% pure. How many ounces of the 60% pure gold must be used to make 100 ounces of an alloy which will be 66% gold?
------------
Equation:
gold + gold = gold
0.70x + 0.60(100-x) = 0.66*100
---
Multiply thru by 100 to get:
70x + 60*100 - 60x = 66*100
10x = 6*100
x = 60 oz (amt. of 70% alloy needed in the mix)
---
100-x = 40 oz (amt of 60% alloy needed in the mix)
==============
Cheers,
Stan H.